OpenStudy (anonymous):

Amanda is decorating her new home and wants to buy some house plants. She is interested in three types of plants costing $7, $10 and $13. If she has budgeted exactly $150 for the plants and wants to buy exactly 15 of them, what are her options?

OpenStudy (anonymous):

You can put this solution on YOUR website! Amanda is decorating her home and wants to buy some house plants. She is interested in three types of plants costing $7, $10, and $13. If she has budgeted $150 for the plants and whats to buy exactly 15 of them, what are her options? let x=number of $7 plants let y=number of $10 plants let z=number of $13 plants .. by trial and error: x=5, y=5, z=5 x+y+z=15 5*7+5*10*5*13=35+50+65=$150 one of her options: buy 5 of each type of plant

OpenStudy (anonymous):

explain how you got that answer